The lenght of a recetangular playing field if 5 ft less than twice its width.If the perimeter of the playing field is 230 feet, find the lenght and width of the field.
Answer by junior403(76)

The lenght of a recetangular playing field if 5 ft less than twice its width.If the perimeter of the playing field is 230 feet, find the lenght and width of the field.
First lets set this problem up as an equation acording to what we already know.
Since we know nothing about the width, lets call that x.
If the length is 5 ft less than twice the width, then The length can be represented as...
2x - 5
The formula for perimeter is 2w + 2l so...
2x + 2(2x - 5) = 230
Now we solve for x...
2x + 4x -10 = 230
6x -10 = 230
Now we can add 10 to both sides of the equation...
6x = 240
Divide by 6...
x = 40
So, the width is 40ft.
To find the length...
2(40)-5 = 75
So, the length is 75ft.
Now we can check our answers to the original equation.
2(40) + 2(75) = 230
80 + 150 = 230
TRUE.
I hope this helps
Good Luck!